# SimSort: A Powerful Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

### v1

@@ 5.1 Benchmarks
BBP L6 dataset (simulation). This benchmark dataset comprised 20 simulated trials, generated as stated in Sec. 4.1, using BBP layer 6 neuron models. It provided 5 ground-truth units per recording for evaluating spike sorting performance.
Hybrid dataset. We prepared this dataset using recordings from Spikeforest (Magland et al., 2020), which combined real neuronal waveforms with background noise. It included 9 static and 9 drift tetrode recordings, each containing 10–15 ground-truth units with a minimum SNR of 3. Details of the dataset preparation are provided in Appendix A.1.
WaveClus dataset. This dataset consisted of four subsets (Easy1, Easy2, Difficult1, and Difficult2), varying noise levels from 0.05 to 0.4, with 3 ground-truth units per recording (Martinez et al., 2009). The original recordings were single-channel data of real neuronal waveforms mixed with background noise. To adapt them for tetrode analysis and increase the overall difficulty, we processed the recordings as described in the Appendix A.1.
IBL Neuropixels dataset. This dataset was derived from the CortexLab KS046 Neuropixels recordings released by International Brain Laboratory (IBL). To compare spike identification performance with CEED (Vishnubhotla et al., 2023), we followed the CEED paper’s method, extracting waveforms from 400 units. Fifty test sets were generated with random seeds, each containing 10 units.
Evaluation Metrics. To evaluate the performance of the overall spike sorting and spike detection tasks, we used Accuracy, Recall, and Precision as the evaluation metrics. Note that these metrics are specifically defined for the spike sorting context, where they measure the alignment between detected spikes and ground truth spike events, rather than their conventional definitions in classification tasks. For the spike identification task, we used the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985) to measure clustering accuracy relative to the ground truth.
@@
5.2 Spike Sorting Results
Simulated dataset
We first examined Simsort on a simulated test set (BBP L6). In Table 1, we evaluated SimSort alongside widely used algorithms, including KiloSort (Pachitariu et al., 2016), KiloSort2, MountainSort4 (Chung et al., 2017), and MountainSort5, implemented via the SpikeInterface pipeline (Buccino et al., 2020), on the simulated BBP L6 dataset. SimSort achieved the best performance across all metrics, demonstrating its robustness and strong zero-shot generalization ability under idealized, simulated conditions.
Table 1: Spike sorting results on BBP L6 dataset. The results for other methods were obtained through SpikeInterface (Buccino et al., 2020). Values are presented as mean $\pm$ S.E.M.
@@ 5.2 Spike Sorting Results
Table 2: Spike sorting results on Hybrid dataset. Data for other methods were sourced from SpikeForest (Magland et al., 2020). Values are mean $\pm$ S.E.M. Best results are in bold. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@
Table 3: Spike sorting results on WaveClus dataset. The results for other methods were obtained through SpikeInterface (Buccino et al., 2020). Values are mean $\pm$ S.E.M. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@
Real-world data-based dataset
@@
Given SimSort’s strong performance on simulated data, we extended our evaluation to real-world data-based datasets to further test its generalizability and robustness (RQ1). On the Hybrid dataset, we compared SimSort against several algorithms via SpikeForest, including HerdingSpikes2 (Hilgen et al., 2017), IronClust, JRClust, KiloSort (Pachitariu et al., 2016), KiloSort2, MountainSort4 (Chung et al., 2017), SpykingCircus, and Tridesclous. The results are summarized in Table 2 and Fig. S1. On this dataset, SimSort demonstrated the highest Accuracy and Precision in static recordings and achieved the best Accuracy and Recall under drift conditions, although it slightly trailed KiloSort in Precision.
@@
For the WaveClus dataset, we used the SpikeInterface pipeline to run KiloSort, KiloSort2, MountainSort4, and MountainSort5. Results are provided in Table 3 and Fig. S1. SimSort performed reliably across both easy and difficult subsets, showing a pronounced advantage on the challenging “difficult” set. This performance was likely due to its ability to handle the unique challenges of this subset, such as lower SNRs and reduced inter-class waveform variability, which resulted in significant overlaps between units. These results underscored SimSort’s effectiveness in tackling noisy and complex scenarios.
@@ 5.3 Detailed Analysis
Table 4: Spike detection results on Hybrid dataset. Values are mean $\pm$ S.E.M. The threshold detector used the best-performing threshold by grid search.
@@
Table 5: Spike identification on Hybrid datasets. ARI values are presented as mean $\pm$ standard deviation. All results were averaged across 20 repeats of GMM.
Methods Hybrid-static Hybrid-drift
PCA 0.79 $\pm$ 0.04 0.74 $\pm$ 0.04
$t$-SNE 0.88 $\pm$ 0.02 0.85 $\pm$ 0.02
UMAP 0.88 $\pm$ 0.02 0.85 $\pm$ 0.01
Ours 0.91 $\pm$ 0.02 0.89 $\pm$ 0.03
@@
After detecting spikes, our spike identification model extracted waveform representations in the latent space, enabling subsequent clustering. To effectively demonstrate the superiority of our identification model, we assembled a waveform dataset by gathering waveform data from each neuron in every hybrid recording. Using this dataset, we compared the 2D UMAP embedding of our model’s inferred representations with features extracted directly from waveforms using widely adopted dimensionality reduction techniques, including PCA (Pearson, 1901), $\boldsymbol{t}$-SNE (van der Maaten & Hinton, 2008), and UMAP (McInnes et al., 2020b). For evaluation, we applied a parametric clustering method, GMM (Dempster et al., 1977), to compute the Adjusted Rand Index (ARI), providing a quantitative measure of the reliability of the representations.
Figure 5: Visualized results of spike identification on an example recording in the Hybrid static subset. The first column illustrated UMAP embeddings of SimSort-learned representations, with GMM clustering results (top) and true labels (bottom). In the second, third, and fourth columns, low-dimensional features were extracted directly from waveform data using UMAP, $t$-SNE, and PCA, respectively. For each method, the top row presented the GMM clustering results, and the bottom row showed the corresponding true labels. See Appendix C for visualization results for additional samples.
@@
As shown in Table 5, SimSort substantially outperformed these conventional methods under both static and drift conditions, achieving the highest average ARI values of 0.91 and 0.89, respectively. These results revealed the robustness and effectiveness of our approach in deriving meaningful and reliable representations from extracellular data.
@@
Additionally, we compared SimSort with CEED, a recent framework for representation learning on extracellular data, using the provided checkpoint and hyperparameters. Both methods were evaluated on fifty test sets from the IBL Neuropixels dataset. As presented in Table S4, Fig. S2, and Fig. S3, our method achieved performance comparable to CEED, with a slightly higher score.
Case Studies
@@
In Fig. 5, we visualized the results of spike identification on an example recording in the Hybrid static subset. SimSort’s embeddings (first column) demonstrated superior clustering performance, with a GMM clustering ARI score of 0.90, closely aligning with true labels. In contrast, UMAP, $t$-SNE, and PCA embeddings (second to fourth columns) yielded lower clustering scores of 0.80, 0.78, and 0.67, respectively. SimSort embeddings showed clear separation between neuron clusters, while other methods exhibited significant overlap, highlighting the advantages of SimSort in learning robust and discriminative features.
Impact of Dimensionality Reduction and Clustering Methods
@@
To evaluate the adaptability of SimSort, we conducted a study to assess the impact of different dimensionality reduction techniques (PCA, $t$-SNE, UMAP, and None) and clustering algorithms (GMM, KMeans (MacQueen, 1967), Mean Shift (Cheng, 1995), and Spectral Clustering (Ng et al., 2001)). Dimensionality reduction was applied to latent features extracted from waveforms, with an additional setup where clustering was performed directly on the raw latent features without dimensionality reduction (None). As illustrated in Fig. 6, the combination of UMAP and Mean Shift achieved the highest accuracy on both hybrid static and drift subsets. Across many other combinations, SimSort also maintained competitive performance, demonstrating its flexibility and robustness to various configurations.
Figure 6: Impact of dimensionality reduction and clustering algorithms on the hybrid dataset. MS denotes Mean Shift.
@@ 5.4 Scaling Law of Data Size
As we emphasized the importance of a large-scale dataset throughout this work, it was interesting to evaluate how SimSort’s performance scaled with the size of the training dataset (RQ3). Specifically, we evaluated performance across sorting accuracy, detection accuracy, and identification ARI score, capturing how each metric varied with the growing size of the training dataset, which ranged from $2^{8}$ to $2^{13}$ trials. As Fig. 7 shows, larger datasets led to almost consistent performance improvements on both hybrid static and drift subsets, reflecting the scaling law of data size.
Figure 7: Performance scaling with increasing dataset size (x-axis). Purple solid lines denoted sorting accuracy, purple dashed lines represented detection accuracy, and green solid lines corresponded to identification ARI. Darker lines indicated the hybrid static subset, while lighter lines corresponded to the drift subset.
@@ 6 Conclusion
In this paper, we presented SimSort, a data-driven framework for automated spike sorting, leveraging a biologically realistic large-scale extracellular dataset. SimSort outperformed commonly used spike sorting algorithms on both simulated and real-world benchmarks, demonstrating its robustness and scalability. Evaluations further highlighted its strong zero-shot transfer capability without fine-tuning.
SimSort has limitations. The models were trained on 4-channel tetrode data. While SimSort shows promising performance on tasks with high-density probes (IBL Neuropixels), future work should extend the training with diverse electrode geometries. Moreover, SimSort only used simulated spike sorting data for training since we focused on validating its zero-shot generalizability. The performance of SimSort can be further improved by incorporating real-world datasets into training.
Impact Statement
Our research contributes to the growing field of neuroscience by providing a promising approach for spike sorting tasks. It leverages pretraining on simulated datasets to enhance scalability and accuracy, offering exciting prospects for future advancements in this domain. We have carefully considered the ethical implications and societal impact of our work and believe it posed no foreseeable risks in these areas.
Acknowledgments
@@ A.1 Data Preparation
Test datasets
Hybrid dataset. This dataset was composed of recordings provided by Spikeforest (Magland et al., 2020). These recordings were generated by Kilosort, with waveform templates recorded at a 5 µm electrode spacing and  1/f background noise. It consisted of two subsets: 9 static tetrode recordings and 9 drift tetrode recordings that simulated sinusoidal probe movements. For each recording, we selected units with an SNR greater than 3 (unit IDs meeting this criterion were provided by SpikeForest to ensure a fair comparison), comprising 10–15 ground-truth units per recording. The dataset we prepared supported two modes of use. To evaluate overall spike sorting and detection performance, we used the raw data as input. For assessing identification performance, we constructed a waveform dataset by extracting 500 waveforms (60-timesteps length) per unit.
@@
A.2 Spike Detection

### v2

@@ 5.1 Benchmarks
Since the models of SimSort has been trained on our simulated dataset, readers must be curious about how well it can generalize to real-world spike sorting tasks. To empirically investigate this problem, we respectively evaluate SimSort in the following benchmarks:
  • Simulated test dataset (with ground truth): BBP L6 dataset from our simulation (using different neuronal models from the training dataset).
  • Real-world data-based synthetic datasets (with label): Hybrid [41], WaveClus [42], IBL Neuropixels [21]. We evaluate both zero-shot (using only our simulated dataset) and fine-tuned (using also the dataset) performance of SimSort.
  • Real-world neural recording (w/o ground truth, indirect validation): extracellular recordings collected from the mouse primary visual cortex during drifting grating stimulation.
@@
Evaluation Metrics. To evaluate the performance of the overall spike sorting and spike detection tasks, we used Accuracy, Recall, and Precision as the evaluation metrics. Note that these metrics are specifically defined for the spike sorting context, where they measure the alignment between detected spikes and ground truth spike events, rather than their conventional definitions in classification tasks. For the spike identification task, we used the Adjusted Rand Index (ARI) [43] to measure clustering accuracy relative to the ground truth.
Table 1: Spike sorting results on BBP L6 dataset. The results for other methods were obtained through SpikeInterface [44]. Values are as mean $\pm$ S.E.M.
@@
5.2 Spike Sorting Results on simulated dataset
We first examined Simsort on a simulated test set (BBP L6). In Table 1, we evaluated SimSort alongside widely used algorithms, including KiloSort [17], KiloSort2, MountainSort4 [6], and MountainSort5, implemented via the SpikeInterface pipeline [44], on the simulated BBP L6 dataset. SimSort achieved the best performance across all metrics, demonstrating its robustness and strong zero-shot generalization ability under idealized, simulated conditions.
Table 2: Spike sorting results on Hybrid dataset. Data for other methods were sourced from SpikeForest [41]. Values are mean $\pm$ S.E.M. Best results are in bold. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@ 5.2 Spike Sorting Results on simulated dataset
Table 3: Spike sorting results on WaveClus dataset. The results for other methods were obtained through SpikeInterface [44]. Values are mean $\pm$ S.E.M. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
@@
5.3 Spiking Sorting results on real-world data-based synthetic datasets
zero-shot generalization
@@ 5.3 Spiking Sorting results on real-world data-based synthetic datasets
Given SimSort’s strong performance on simulated data, we extended our evaluation to real-world data-based datasets to further test its zero-shot generalizability and robustness (RQ1). On the Hybrid dataset, we compared SimSort against several algorithms via SpikeForest, including HerdingSpikes2 [18], IronClust, JRClust, KiloSort [17], KiloSort2, MountainSort4 [6], SpykingCircus, and Tridesclous. The results are summarized in Table 2 and Fig. S1. On this dataset, SimSort demonstrated the highest Accuracy and Precision in static recordings and achieved the best Accuracy and Recall under drift conditions, although it slightly trailed KiloSort in Precision.
@@
For WaveClus dataset, we used SpikeInterface pipeline to run KiloSort, KiloSort2, MountainSort4, and MountainSort5. Results are provided in Table 3 and Fig. S1. SimSort performed reliably across both easy and difficult subsets, showing a pronounced advantage on the challenging “difficult” set. This performance was likely due to its ability to handle the unique challenges of this subset, such as lower SNRs and reduced inter-class waveform variability, which resulted in significant overlaps between units. These results underscored SimSort’s effectiveness in noisy and complex scenarios.
Fine-tuning
Beyond demonstrating SimSort’s zero-shot transfer ability on unseen datasets, we also investigate its performance after fine-tuning on a small amount of data from the target domain (Table 4). While the overall improvements in sorting metrics are modest, we observe notable changes in spike detection performance, particularly in the balance between Precision and Recall. These changes suggest that fine-tuning helps the model adapt to dataset-specific waveform and noise characteristics, enhancing its practicality with minimal labeled data.
Table 4: Performance metrics of SimSort after fine-tuning on different datasets.The hybrid-static and hybrid-drift datasets were fine-tuned on recordings 1-3 and tested on recordings 4-9 (train/test set split). The waveclus -easy and -difficult datasets were fine-tuned on subsets easy1/diff1 and tested on easy2/diff2, respectively. The IBL Neuropixels dataset was fine-tuned on 100 neurons and evaluated on 50 test sets, each with 10 randomly selected neurons.
Dataset Model Spike Detection (Acc/Rec/Prec) Spike Identification (ARI) Spike Sorting (Acc/Rec/Prec)
Hybrid-static SimSort 0.71±0.02 / 0.83±0.02 / 0.83±0.02 0.90±0.02 0.59±0.03 / 0.65±0.03 / 0.76±0.02
SimSort-FT 0.75±0.03 / 0.89±0.02 / 0.82±0.02 0.92±0.02 0.58±0.03 / 0.65±0.03 / 0.73±0.02
Hybrid-drift SimSort 0.67±0.02 / 0.80±0.02 / 0.81±0.01 0.88±0.03 0.54±0.09 / 0.61±0.09 / 0.67±0.07
SimSort-FT 0.73±0.03 / 0.85±0.03 / 0.83±0.01 0.88±0.02 0.56±0.04 / 0.63±0.03 / 0.69±0.03
Waveclus-easy SimSort 0.59±0.02 / 0.99±0.00 / 0.59±0.02 0.91±0.01 0.70±0.11 / 0.79±0.07 / 0.84±0.09
SimSor-FT 0.94±0.01 / 0.96±0.00 / 0.98±0.01 0.91±0.01 0.73±0.08 / 0.79±0.05 / 0.91±0.04
Waveclus-difficult SimSort 0.66±0.08 / 0.99±0.00 / 0.67±0.08 0.88±0.01 0.76±0.14 / 0.84±0.09 / 0.84±0.12
SimSort-FT 0.95±0.01 / 0.96±0.00 / 0.99±0.01 0.87±0.03 0.77±0.14 / 0.82±0.10 / 0.86±0.13
IBL Neuropixels SimSort N/A 0.49±0.09 N/A
SimSort-FT N/A 0.53±0.09 N/A
5.4 Real-World Neural Recording
Beyond quantitative benchmarks, we further assessed SimSort in a real experimental setting. Specifically, we applied it to extracellular recordings (collected in our wet lab, see Appendix A.1.2) from the mouse primary visual cortex obtained using tetrodes during presentation of full-field drifting gratings in 8 orientations (Fig. 5a). From a representative recording session, SimSort identified six putative single units (Fig. 5b). These units exhibited distinct spike waveforms and their autocorrelograms showed clear refractory period structure(Fig. 5c), suggesting good isolation. We further analyzed the orientation tuning of these units. Several exhibited selective responses, with global orientation selectivity index (gOSI) values exceeding 0.2 (Fig. 5d; Appendix A for definition). Across 10 tetrode recordings, SimSort isolated a total of 40 units, with a mean gOSI of 0.27 ± 0.14 (mean ± S.D., Fig. S17). These results demonstrate the practical applicability of SimSort to real neural recordings.
Figure 5: Evaluation of SimSort on real experimental data (See Fig. S17 for additional recording results). (a) Experimental setup: extracellular recordings from the mouse primary visual cortex during drifting gratings in eight directions. (b) Visualization of SimSort’s sorting results. (c) Autocorrelograms and average waveform of each unit. (d) Orientation tuning curves and global orientation selectivity index (gOSI) for each unit.
@@ 5.5 Detailed Analysis
Table 5: Spike detection and identification results on the Hybrid dataset. Values are mean $\pm$ S.E.M.
(a) Spike detection results. The threshold detector used the best-performing threshold by grid search.
@@
(b) Spike identification results. All results averaged across 20 repeats of GMM.
Dataset PCA $t$-SNE UMAP Ours
Hybrid-static 0.79 $\pm$ 0.04 0.88 $\pm$ 0.02 0.88 $\pm$ 0.02 0.91 $\pm$ 0.02
Hybrid-drift 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02 0.85 $\pm$ 0.01 0.89 $\pm$ 0.03
@@
After detecting spikes, our spike identification model extracted waveform representations in the latent space, enabling subsequent clustering. To effectively demonstrate the superiority of our identification model, we assembled a waveform dataset by gathering waveform data from each neuron in every hybrid recording. Using this dataset, we compared the 2D UMAP embedding of our model’s inferred representations with features extracted directly from waveforms using widely adopted dimensionality reduction techniques, including PCA [23], $\boldsymbol{t}$-SNE [24], and UMAP [38]. For evaluation, we applied a parametric clustering method, GMM [39], to compute the Adjusted Rand Index (ARI), providing a quantitative measure of the reliability of the representations.
@@
As shown in Table 5(b), SimSort substantially outperformed these conventional methods under both static and drift conditions, achieving the highest average ARI values of 0.91 and 0.89, respectively. These results revealed the robustness and effectiveness of our approach in deriving meaningful and reliable representations from extracellular data.
@@
Additionally, we compared SimSort with CEED [21], a recent framework for representation learning on extracellular data, using the author-provided model. Note that CEED is trained on datasets derived from IBL DanLab DY016 and DY009 Neuropixels recordings, which are more closely related to our test sets. We evaluated both SimSort and CEED on fifty test sets from the IBL Neuropixels dataset (specifically derived from the IBL CortexLab KS046 recording). As presented in Table S4, Fig. S2S3, our method, without training on the dataset, achieved performance comparable to CEED.
@@
Case Studies: In Fig. 6, we visualized the results of spike identification on an example recording in the Hybrid static subset. SimSort’s embeddings (first column) demonstrated superior clustering performance, with a GMM clustering ARI score of 0.90, closely aligning with true labels. In contrast, UMAP, $t$-SNE, and PCA embeddings (second to fourth columns) yielded lower clustering scores of 0.80, 0.78, and 0.67, respectively. SimSort embeddings showed clear separation between neuron clusters, while other methods exhibited significant overlap, highlighting the advantages of SimSort in learning robust and discriminative features.
@@
Impact of Dimensionality Reduction and Clustering Methods: To evaluate the adaptability of SimSort, we conducted a study to assess the impact of different dimensionality reduction techniques (PCA, $t$-SNE, UMAP, and None) and clustering algorithms (GMM, KMeans [45], Mean Shift [40], and Spectral Clustering [46]). Dimensionality reduction was applied to latent features extracted from waveforms, with an additional setup where clustering was performed directly on the raw latent features without dimensionality reduction (None). As illustrated in Fig. S4, the combination of UMAP and Mean Shift achieved the highest accuracy on both hybrid static and drift subsets. Across many other combinations, SimSort also maintained competitive performance, demonstrating its flexibility and robustness to various configurations.
Figure 6: Visualized results of spike identification on an example recording in the Hybrid static subset. See Appendix C for visualization results for additional samples.
@@ 5.6 Scaling Law of Data Size
Figure 7: Performance scaling with increasing training data size on the Hybrid-drift and -static.
As we emphasized the importance of a large-scale dataset throughout this work, it was interesting to evaluate how SimSort’s performance scaled with the size of the training dataset (RQ3). Specifically, we evaluated performance across sorting accuracy, detection accuracy, and identification ARI score, capturing how each metric varied with the growing size of the training dataset, which ranged from $2^{8}$ to $2^{13}$ trials. As Fig. 7 shows, larger datasets led to almost consistent performance improvements on both hybrid static and drift subsets, reflecting that the large-scale dataset holds the key of SimSort’s effectiveness.
@@ 6 Conclusion
In this paper, we presented SimSort, a data-driven framework for automated spike sorting, built upon a biologically realistic large-scale extracellular dataset. SimSort achieves competitive or superior performance compared to commonly used spike sorting algorithms on both simulated and real-world benchmarks, demonstrating its robustness and scalability. Evaluations further highlighted its zero-shot transfer capability.
SimSort has limitations. The models were trained on 4-channel tetrode data. While SimSort shows promising performance on tasks with high-density probes (IBL Neuropixels), future work should extend the training with diverse electrode geometries.
Acknowledgments and Disclosure of Funding
@@ A.1.1 Train datasets
A.1.2 Test datasets
BBP L6 dataset (simulation). This dataset comprised 20 simulated trials, generated as stated in Sec. 4.1, using BBP layer 6 neuron models. It provided 5 ground-truth units per recording for evaluating spike sorting performance.
Hybrid dataset. This dataset was composed of recordings provided by Spikeforest [41]. These recordings were generated by Kilosort, with waveform templates recorded at a 5 µm electrode spacing and  1/f background noise. It consisted of two subsets: 9 static tetrode recordings and 9 drift tetrode recordings that simulated sinusoidal probe movements. For each recording, we selected units with an SNR greater than 3 (unit IDs meeting this criterion were provided by SpikeForest to ensure a fair comparison), comprising 10–15 ground-truth units per recording. The dataset we prepared supported two modes of use. To evaluate overall spike sorting and detection performance, we used the raw data as input. For assessing identification performance, we constructed a waveform dataset by extracting 500 waveforms (60-timesteps length) per unit.
@@ A.1.2 Test datasets
Real-world Neural Recording. These recordings were collected by our wet lab (approved by research ethics committee). A 16-channel tetrode was implanted unilaterally in the primary visual cortex (V1) of mice. During the recording sessions, the mice were head-fixed, and the contralateral eye was stimulated with drifting grating stimuli of 8 different orientations. The stimuli were presented in a randomized sequence over 8 repeated trials. Each grating had a spatial frequency of 0.04 cycles per degree (cpd) and a temporal frequency of 2 Hz, lasting for 1.5 seconds per trial.
To quantify the orientation tuning of individual neurons, we computed the global orientation selectivity index (gOSI), as defined in the equation below:
$gOSI=\frac{\left\|\sum R(\theta)e^{2i\theta}\right\|}{\sum R(\theta)}$
where $\theta$ is the direction of the moving grating and $R(\theta)$ is the neuronal response to direction $\theta$. The imaginary unit is denoted as $i$ whose square is $-1$. The gOSI ranges from 0 to 1, with 0 indicating no orientation selectivity (equal response to all directions), and 1 indicating maximal selectivity (response only to a single orientation).
@@
Real-world Neural recording 10 40000 4 $\approx 250$ / No Ground-truth
A.2 Technical Details of Simulation
Intracellular Simulation. Intracellular electrophysiological activity was simulated using the NEURON package [30, 47]. Pink noise, scaled to the rheobase (the minimum current needed to elicit an action potential in a neuron), was injected into each neuron’s soma to evoke stochastic action potentials, simulating biologically realistic background noise. The temporal resolution was set to 0.1 ms, and simulations spanned a total duration of 600 seconds. Neurons were initialized with a resting membrane potential of -70 mV and maintained at a physiological temperature of 34°C. Synaptic mechanisms and biophysical properties were dynamically compiled and loaded to ensure compatibility with the NEURON simulation environment.
For each neuron, a multi-compartmental model was used to calculate transmembrane currents [48, 49, 50, 51]. The neuron was divided into compartments, each described as an equivalent electrical circuit. The dynamics of the membrane potential in each compartment $n$ were governed by Kirchhoff’s current law [52], where the sum of all currents entering or leaving a node must equal zero. Considering that extracellular potential changed much slower than ion channel dynamics, we assumed it constant. The temporal evolution of the membrane potential $V_{n}$ in compartment $n$ was given by $g_{n,n+1}(V_{n+1}-V_{n})-g_{n-1,n}(V_{n}-V_{n-1})=C_{n}\frac{dV_{n}}{dt}+\sum_{j}I_{n}^{(j)},$ where $g_{n,n+1}$ and $g_{n-1,n}$ represented the conductances between neighboring compartments, $C_{n}$ was the membrane capacitance, and $\sum_{j}I_{n}^{(j)}$ accounted for ionic currents and any externally applied currents.
Extracellular Simulation. Extracellular potentials were modeled using the volume conductor theory, which links transmembrane currents to extracellular potentials recorded at electrode sites [53, 54, 55]. In each simulation trial, five neurons were randomly selected from the available models and positioned within a 100 × 100 × 100 $\mu$m3 cubic space. These neurons consisted of different types to represent the diversity of cortical microcircuits. A virtual tetrode electrode with four recording sites was placed randomly within the same space to capture extracellular potentials.
A.3 Spike Detection